OpenStudy (ghosting):

How do I do this one? A standard deck of cards has 4 suits (hearts, diamonds, spades, and clubs) each with 13 cards in it (ace -10, jack, queen, and king). The number cards are 2-10. When drawing a card randomly from a deck, what is the probability that it will be either a diamond or a number card? Type your answer as a fraction reduced to the simplest form, like this: 7/11

OpenStudy (misty1212):

there are 13 diamonds and 3*9=27 other "number cards" for a total of 13 + 27 = 40

OpenStudy (misty1212):

so reduce \[\frac{40}{52}\]

OpenStudy (ghosting):

10/13!!!